What is the resultant velocity of a balloon that rises at a rate of 5 m/s and has the wind blowing it parallel to the ground at 3 m/s?
Answer #1

Solution)

Velocity in y-axis=5 m/s

Vekocity in x-axis=3 m/s

So, resultant velocity= sqrt(5^2 + 3^2)= sqrt(34)=5.83 m/s

============